# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ VIII Acknowledgments
We would like to thank Matthew Jones, Chris Laumann, Gianni Mossi, Eleanor Rieffel, Antonello Scardicchio and Davide Venturelli for valuable discussions of the issues in this work. We would like to thank Caleb Rotello for detailed feedback on the manuscript. We would also like to thank Takuto Komatsuki and Joey Liu for support with HPC calculations. This work was supported by the DARPA Reversible Quantum Machine Learning and Simulation program under contract HR00112190068, as well as by National Science Foundation grants PHY-1653820, PHY-2210566, DGE-2125899, and by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. The SQMS Center supported EK’s advisory role in this project, as well as time improving and fine tuning the algorithms and writing the paper. Many of the numerical simulations in this work were performed with a generous grant of HPC access from the Fujitsu Corporation. Part of this research was performed while the one of the authors (BAB) was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). The Flatiron Institute is a division of the Simons Foundation.
Appendix A Tunneling between two $p$-spin wells
@@
While fully sufficient for our purposes here (where $p=3$), we want to highlight two issues with the formulation in Eq. IV.3 that suggest a more refined treatment should be developed to tackle $p>3$ and/or asymmetric minima. First, $\partial_{m}E_{m,n}$ is guaranteed to vanish at some point, since the energy curve at the peak of the barrier is smooth, and since it enters the calculation in the denominator, it predicts a diverging scale for energy corrections. In symmetric cases, this divergence is canceled by the factor of $N/2-2n$ in the numerator and the result is finite, but if the energy curve is not symmetrical between the two minima those two factors won’t generally coincide (as the barrier peak won’t occur at $M/2$ flips for asymmetric minima) and the divergence will not be canceled.
@@
Figure 10: Comparison of Eq. IV.3 to the tunnel splitting found from exact diagonalization of the Hamiltonian in Eq. IV.3, for $p=\left\{2,3,4,5,7,9\right\}$ (blue, gold, green, purple, red, brown) and $\kappa$ stated in the text. There are no free parameters in this; exponents found by extrapolating Eq. IV.3 out to $N=80$ agree with the numerical results to within few percent. To make sensible predictions for $p\geq 4$ an additional modification was needed to regularize the transverse field corrections to state energies– see text for details. Only $p=3$ is directly relevant to this work; simulation of other $p$ values is to demonstrate the accuracy and flexibility of our theory.
One can also work through this analysis for the splitting $p=2$ mean field all-to-all ferromagnet, where the transition out of the ferromagnet phase occurs when $\kappa=2$ for this normalization. In that case all $N$ spins must flip and there are no secondary spin corrections; further, due to a symmetry cancellation the self energy corrections (e.g. Eq. IV.3) are energy-independent and thus do not renormalize the tunneling barrier. Repeating the same steps for this somewhat simpler calculation yields
@@
Of note is that Eq. 37 is both of similar quantitative accuracy as the results plotted in FIG. 10, and it predicts that the decay exponent smoothly approaches zero as $\kappa\to 2$, at which point the minimum gap decays inverse polynomially. Similar behavior–a crossover to polynomial scaling when theory predicts the decay exponent must vanish–was observed both analytically and numerically in the 1d transverse field Ising chain in Grattan et al. (2023). We see this as lending further indirect support to our polynomial time approximation hardness prediction in Eq. IV.4–though the situation is fairly different all these calculations identify a crossover to polynomial scaling by the point at which a predicted decay exponent vanishes in careful many-body theory.
Appendix B Band structure considerations for linear spectral folding
@@
Finally, let $\Omega_{Lj}$ be the $M$-spin tunneling matrix element into a bitstring state $\left|T_{j}\right>$ from $\left|L_{D}\right>$, where $\Omega_{Lj}\right>_{j}=\Omega_{0}$ as calculated in Eq. IV.3.
@@
Thus, the average squared matrix element to tunnel into a state in the band center has the same scaling as one obtains from a more naive calculation that ignores the band structure entirely. Since the majority of states in the band are near the center in this respect, this calculation shows that ignoring the local band structure in $\mathcal{T}$ is a decent approximation for obtaining the average collective tunneling rate that enters into equations like IV.4.
But what happens when we consider states that are more extremal, e.g. near the top or bottom of the band? Near the top of the band, we can assume significant local alternation in the signs/phases of the $c_{jk}$ coefficients, and thus the assumption of randomization is still fairly good and Eq. B likely accurately captures the scaling. For states near the bottom of the band, however, the situation changes. First, these states will be lower in energy by an $O\left(N\right)$ factor, which means $C\left(t\right)$ must be further reduced for $\left|L_{D}\right>$ to cross them, and the individual tunneling matrix elements $\Omega_{Lj}$ into such states may scale differently as a result. We label these matrix elements $\Omega_{Lj}^{\prime}$ and do not attempt to predict what, if any, changes to scaling may arise.
Second, to minimize the energy the signs/phases of the $c_{jk}$ of nearby (in Hamming distance) configurations will be synchronized given that the matrix elements that couple them are real and (mostly) negative (again, in this gauge choice). And this synchronization can dramatically enhance tunneling rates. Let us guess, for example, that in the band ground state $\left|\psi_{0}\right>$, most $c_{j0}$ are positive. Then:
@@
Figure 11: Probabilities of returning states within Hamming distance $D_{H}=\left\{2/5,1/3,1/4,1/5,1/8,0\right\}N$ flips from the ground state $G$ for QAOA (top to bottom curves), for $N_{C}=\left\{2,4,6,3\sqrt{N}/2\right\}N$ (clockwise from top left) and application parameters in text. As seen in the figures, the per-shot probability of finding the ground state is essentially independent of constraint density and tracks the prediction $N2^{-0.28N}$ in Eq. 15. On the other hand, the probabilities of returning states at various extensive fractional Hamming distances, e.g. $N/4$ or fewer flips, decay much more slowly, and the task of finding states comparatively close to $G$ becomes easier as $N_{C}/N$ increases. That said, these probabilities all decay exponentially with $N$ in all cases, consistent with the worsening approximation ratios observed in figure 6.

### v3

@@ VIII Acknowledgments
We would like to thank Tameem Albash, David Huse, Matthew Jones, Chris Laumann, Gianni Mossi, Ojas Parekh, Eleanor Rieffel, Luca Trevisan, Antonello Scardicchio and Davide Venturelli for valuable discussions of the issues in this work. We would like to thank Alex Dalzell and Caleb Rotello for detailed feedback on the manuscript. We would also like to thank Takuto Komatsuki and Joey Liu for support with HPC calculations. This work was supported by the DARPA Reversible Quantum Machine Learning and Simulation program under contract HR00112190068, as well as by National Science Foundation grants PHY-1653820, PHY-2210566, DGE-2125899, and by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359. The SQMS Center supported EK’s advisory role in this project, as well as time improving and fine tuning the algorithms (including the formulation of the MSFO algorithm) and writing the paper. Many of the numerical simulations in this work were performed with a generous grant of HPC access from the Fujitsu Corporation. Part of this research was performed while the one of the authors (BAB) was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). The Flatiron Institute is a division of the Simons Foundation.
Appendix A Appendix: small-$N$ deviations from the asymptotic cost per flip curve
In this section, we consider small-$N$ deviations from the asymptotic mean cost per flip scaling, Eq. 18. As discussed in that subsection, the statement $E_{\mathrm{avg}}\left(x\right)=E\left(G\right)\left(1-\frac{2x}{N}\right)^{p}$ is only asymptotically exact, owing to the approximation made in taking the step size to zero in Eq. 17. Empirically, as plotted in FIG. 14, the real average energy at finite $N$ tends to be larger (e.g. less negative) than the asymptotic scaling predicts, though as fraction of the total energy these deviations decrease with system size, with an overall envelope that scales roughly as $1/N$.
While this effect has no impact on the asymptotic scaling of our methods, it can potentially pose challenges for small $N$ simulation. As the average energy costs (e.g. $E_{\mathrm{avg}}\left(m\right)-E\left(G\right)$) enter in products into the denominators of collective tunneling rate calculations, even small fractional increases can produce significant reductions in the associated minimum gaps. They can similarly alter how the system responds to the spectral filtering transformations. However, this is at least partially balanced in scaling by the fact that the scale of the cost per flip increases is decreasing with $N$, so for the algorithms considered in this work the resulting deviations from the predicted asymptotic scaling are typically modest.
We caution however that at larger $p$ these deviations become more severe, even if they too vanish as $N\to\infty$. This issue is one of the main reasons we focused exclusively on $p=3$ MAX-3-XORSAT in this work, as the methods and techniques we introduced can readily generalize to higher degree constraints. We feel that the generalization from mean-field like analytical calculations to real disordered problems is highly nontrivial, and thought it important to be able to test the basic veracity of our predictions using detailed numerical simulations. For $p=5$ we were not confident in being able to access the asymptotic scaling of the system in full wavefunction simulations with $N\leq 25$, so we did not explore MAX-5-XORSAT in this work.
Figure 14: Fractional deviations from the average energy Eq. 18 for random MAX-3-XORSAT instances, as a function of $m$ flips away from the planted ground state $G$, numerically estimated for $N=\left\{12,24,36,48,60\right\}$ (blue circles through purple triangles and monotonically decreasing).
Appendix B Tunneling between two $p$-spin wells
To benchmark our prediction in Eq. IV.3.1, we performed a series of numerical simulations, shown in FIG. 15. For these calculations, we computed the degeneracy splitting $2\Omega_{0}$ for two p-spin wells centered at classical states $a$ and $b$ separated $N/2$ flips apart, with a total Hamiltonian given by
$-\kappa\sum_{j}X_{j},$ (47)
@@ Appendix B Tunneling between two $p$-spin wells
While fully sufficient for our purposes here (where $p=3$), we want to highlight two issues with the formulation in Eq. IV.3.1 that suggest a more refined treatment should be developed to tackle $p>3$ and/or asymmetric minima. First, $\partial_{m}E_{m,n}$ is guaranteed to vanish at some point, since the energy curve at the peak of the barrier is smooth, and since it enters the calculation in the denominator, it predicts a diverging scale for energy corrections. In symmetric cases, this divergence is canceled by the factor of $N/2-2n$ in the numerator and the result is finite, but if the energy curve is not symmetrical between the two minima those two factors won’t generally coincide (as the barrier peak won’t occur at $M/2$ flips for asymmetric minima) and the divergence will not be canceled.
@@
Figure 15: Comparison of Eq. IV.3.1 to the tunnel splitting found from exact diagonalization of the Hamiltonian in Eq. B, for $p=\left\{2,3,4,5,7,9\right\}$ (blue, gold, green, purple, red, brown) and $\kappa$ stated in the text. There are no free parameters in this; exponents found by extrapolating Eq. IV.3.1 out to $N=80$ agree with the numerical results to within few percent. To make sensible predictions for $p\geq 4$ an additional modification was needed to regularize the transverse field corrections to state energies– see text for details. Only $p=3$ is directly relevant to this work; simulation of other $p$ values is to demonstrate the accuracy and flexibility of our theory.
One can also work through this analysis for the splitting $p=2$ mean field all-to-all ferromagnet, where the transition out of the ferromagnet phase occurs when $\kappa=2$ for this normalization. In that case all $N$ spins must flip and there are no secondary spin corrections; further, due to a symmetry cancellation the self energy corrections (e.g. Eq. IV.3.1) are energy-independent and thus do not renormalize the tunneling barrier. Repeating the same steps for this somewhat simpler calculation yields
@@
Of note is that Eq. 49 is both of similar quantitative accuracy as the results plotted in FIG. 15, and it predicts that the decay exponent smoothly approaches zero as $\kappa\to 2$, at which point the minimum gap decays inverse polynomially. Similar behavior–a crossover to polynomial scaling when theory predicts the decay exponent must vanish–was observed both analytically and numerically in the 1d transverse field Ising chain in Grattan et al. (2023). We see this as lending further indirect support to our polynomial time approximation hardness prediction in Eq. IV.3.2–though the situation is fairly different all these calculations identify a crossover to polynomial scaling by the point at which a predicted decay exponent vanishes in careful many-body theory.
Appendix C Band structure considerations for linear spectral folding
@@
Finally, let $\Omega_{Lj}$ be the $M$-spin tunneling matrix element into a bitstring state $\left|T_{j}\right>$ from $\left|L_{D}\right>$, where $\Omega_{Lj}\right>_{j}=\Omega_{0}$ as calculated in Eq. IV.3.1.
@@
Thus, the average squared matrix element to tunnel into a state in the band center has the same scaling as one obtains from a more naive calculation that ignores the band structure entirely. Since the majority of states in the band are near the center in this respect, this calculation shows that ignoring the local band structure in $\mathcal{T}$ is a decent approximation for obtaining the average collective tunneling rate that enters into equations like IV.3.2.
But what happens when we consider states that are more extremal, e.g. near the top or bottom of the band? Near the top of the band, we can assume significant local alternation in the signs/phases of the $c_{jk}$ coefficients, and thus the assumption of randomization is still fairly good and Eq. C likely accurately captures the scaling. For states near the bottom of the band, however, the situation changes. First, these states will be lower in energy by an $O\left(N\right)$ factor, which means $C\left(t\right)$ must be further reduced for $\left|L_{D}\right>$ to cross them, and the individual tunneling matrix elements $\Omega_{Lj}$ into such states may scale differently as a result. We label these matrix elements $\Omega_{Lj}^{\prime}$ and do not attempt to predict what, if any, changes to scaling may arise.
Second, to minimize the energy the signs/phases of the $c_{jk}$ of nearby (in Hamming distance) configurations will be synchronized given that the matrix elements that couple them are real and (mostly) negative (again, in this gauge choice; see Bravyi et al. (2006)). And this synchronization can dramatically enhance tunneling rates. Let us guess, for example, that in the band ground state $\left|\psi_{0}\right>$, most $c_{j0}$ are positive. Then: